# Joint multi-band deconvolution for Euclid and Vera C. Rubin images

### v1

@@ Appendix B Plug-and-Play ADMM
Plug-and-Play Alternating Direction Method of Multipliers (PnP ADMM) has emerged as a powerful framework for solving inverse problems by combining iterative optimization techniques with deep learning-based priors. Originally developed for convex optimization problems with linear equality constraints (Boyd et al. 2010), ADMM decomposes the minimization process into sequential sub-problems, typically involving a data fidelity term and a regularization term, followed by an update of the dual variable. Previous works (Venkatakrishnan et al. 2013; Sreehari et al. 2016; H. Chan et al. 2016) have interpreted these sub-steps as an inversion step followed by a denoising step, coupled via the augmented Lagrangian term and the dual variable. The ADMM PnP approach extends this idea by replacing the proximal operator related to the prior with a deep neural network (DNN) trained as a denoiser (Meinhardt et al. 2017; Bigdeli et al. 2017; Gupta et al. 2018; Sureau et al. 2020), allowing for greater flexibility in handling complex image priors. Compared to direct deep-learning-based inverse models, PnP ADMM offers several advantages: (1) it decouples the inversion step from the DNN, enabling the inclusion of additional convex constraints that can be efficiently handled via optimization, (2) it reduces the cost of learning by focusing solely on training a denoiser rather than multiple networks, as seen in unfolding approaches, and (3) by iterating between denoising and inversion, it ensures that the network output remains consistent with the observed data. In this work, we integrate the PnP ADMM framework from Sureau et al. (2020) with the DRUNet denoiser from Zhang et al. (2022). Specifically, we employ DRUNet in the proximal update step of the PnP framework to enhance the denoising performance.
@@ B.1 The proposed solution
Considering the forward model described in Section 2.1, we defined the following loss functions, akin to equations 6-8, but incorporating an additional augmented Lagrangian term with the dual variable. These functions were then minimized using the algorithm outlined in Section B.2.
@@
where $\text{where }\textbf{K}=\left\|\frac{\hskip 2.0pt\mathbf{h}_{euc}\ast\sum\limits_{c\in\{r,i,z\}}\alpha_{c}\mathbf{x}_{c}-\mathbf{y}_{euc}}{\sigma_{euc}}\right\|_{F}^{2}$ (20)
As before, the first terms in equations 17-19 represent the data fidelity terms for each respective band, with $\sigma_{r}$, $\sigma_{i}$, $\sigma_{z}$ being the corresponding noisemaps. The second terms represent the augmented Lagrangian, incorporating the dual variable $\mathbf{z}$ to split the problem into two sub-problems: an inversion/deconvolution step followed by a denoising step. The third terms, as explained in Section 2.2, are the constraining terms that enforce the condition that the sum of images in the Rubin $r$-, $i$-, and $z$-band images equals the $Euclid$ $VIS$-band images. For the ADMM update, the parameter $\rho$ was manually tuned based on the approach from Sureau et al. (2020) to strike a balance between quickly stabilizing the algorithm (with a higher $\rho$) and prioritizing the minimization of the data fidelity term in the early iterations (with a lower $\rho$). The values of all other hyperparameters have been previously described in Sections 2.2 and 2.3.
@@ B.3 Results
The algorithm simultaneously processed the noisy simulations from the three Vera C. Rubin bands and the Euclid $VIS$-band, along with their respective PSFs. These noisy images served as initializations or first guesses. The algorithm was run for 200 iterations, with convergence typically observed within $150$-$200$ iterations for all images in our dataset. Figure 11 shows the convergence plot of the loss function for the deconvolved output in Figure LABEL:subfig:mcdec1_pnp.
Figure 11: The loss function for the galaxy shown in Figure LABEL:subfig:mcdec1_pnp. Convergence is achieved at around 200 iterations when the relative change in loss value $<10^{-3}$ and the curve is flat.
Figure LABEL:fig:deconv_outputs_pnp presents the same two examples of deconvolved images as shown in Figure LABEL:fig:deconv_outputs. Qualitatively, the outputs for the two methods closely resemble each other. Compared to the original algorithm, there is a slight reduction in NMSE by approximately $0.26\%$. However, due to the additional computational steps involved in the iterative process and the proximal denoising step, PnP ADMM takes approximately 50 times longer to run than the original algorithm. This is consistent with the findings of Sureau et al. (2020). Hence, this significant increase in computational time makes it less practical for use on large datasets.

### v2

@@ Appendix B Plug-and-play ADMM
The plug-and-play alternating direction method of multipliers (PnP ADMM) has emerged as a powerful framework for solving inverse problems by combining iterative optimization techniques with deep-learning-based priors. Originally developed for convex optimization problems with linear equality constraints (Boyd et al. 2010), ADMM decomposes the minimization process into sequential sub-problems, typically involving a data fidelity term and a regularization term, followed by an update of the dual variable. Previous works (Venkatakrishnan et al. 2013; Sreehari et al. 2016; H. Chan et al. 2016) have interpreted these sub-steps as an inversion step followed by a denoising step, coupled via the augmented Lagrangian term and the dual variable. The PnP ADMM approach extends this idea by replacing the proximal operator related to the prior with a deep neural network (DNN) trained as a denoiser (Meinhardt et al. 2017; Bigdeli et al. 2017; Gupta et al. 2018; Sureau et al. 2020), allowing for greater flexibility in handling complex image priors. Compared to direct deep-learning-based inverse models, PnP ADMM offers several advantages: (1) it decouples the inversion step from the DNN, enabling the inclusion of additional convex constraints that can be efficiently handled via optimization, (2) it reduces the cost of learning by focusing solely on training a denoiser rather than multiple networks, as seen in unfolding approaches, and (3) by iterating between denoising and inversion, it ensures that the network output remains consistent with the observed data. In this work, we integrate the PnP ADMM framework from Sureau et al. (2020) with the DRUNet denoiser from Zhang et al. (2022). Specifically, we employ DRUNet in the proximal update step of the PnP framework to enhance the denoising performance.
@@ B.1 The proposed solution
Considering the forward model described in Sect. 2.1, we defined the following loss functions, akin to Eqs. 6-8, but incorporating an additional augmented Lagrangian term with the dual variable. These functions were then minimized using the algorithm outlined in Appendix B.2.
@@
where $\text{where }\textbf{K}=\left\|\frac{\hskip 2.0pt\mathbf{h}_{euc}\ast\sum\limits_{c\in\{r,i,z\}}\alpha_{c}\mathbf{x}_{c}-\mathbf{y}_{euc}}{\sigma_{euc}}\right\|_{F}^{2}.$ (20)
As before, the first terms in Eqs. 17-19 represent the data fidelity terms for each respective band, with $\sigma_{r}$, $\sigma_{i}$, and $\sigma_{z}$ being the corresponding noise maps. The second terms represent the augmented Lagrangian, incorporating the dual variable $\mathbf{z}$ to split the problem into two sub-problems: an inversion/deconvolution step followed by a denoising step. The third terms, as explained in Sect. 2.2, are the constraining terms that enforce the condition that the sum of the Rubin $r$-, $i$-, and $z$-band images equals the Euclid VIS-band images. For the ADMM update, the parameter $\rho$ was manually tuned based on the approach from Sureau et al. (2020) to strike a balance between quickly stabilizing the algorithm (with a higher $\rho$) and prioritizing the minimization of the data fidelity term in the early iterations (with a lower $\rho$). The values of all other hyperparameters have been previously described in Sects. 2.2 and 2.3.
@@ B.3 Results
The algorithm simultaneously processed the noisy simulations from the three Rubin bands and the Euclid VIS band, along with their respective PSFs. These noisy images served as initializations or first guesses. The algorithm was run for 200 iterations, with convergence typically observed within $150$-$200$ iterations for all images in our dataset. Figure 11 shows the convergence plot of the loss function for the deconvolved output in Fig. LABEL:subfig:mcdec1_pnp.
Figure 11: Loss function for the galaxy shown in Fig. LABEL:subfig:mcdec1_pnp. Convergence is achieved at around 200 iterations when the relative change in loss value is $<10^{-3}$ and the curve is flat.
Figure LABEL:fig:deconv_outputs_pnp presents the same two examples of deconvolved images as shown in Fig. LABEL:fig:deconv_outputs. Qualitatively, the outputs for the two methods closely resemble each other. Compared to the original algorithm, there is a slight reduction in NMSE by approximately $0.26\%$. However, due to the additional computational steps involved in the iterative process and the proximal denoising step, PnP ADMM takes approximately 50 times longer to run than the original algorithm. This is consistent with the findings of Sureau et al. (2020). Hence, this significant increase in computational time makes it less practical for use on large datasets.